OpenStudy (anonymous):

Let x be the speed of the boat you are taking to the island. The other boats speed can then be written relative to the variable x. Write the speed of each boat using the variable x

OpenStudy (anonymous):

You are traveling to an island by boat. You plan to take one boat there and one boat back. Determine how fast each one travels. Steam boat makes the trip in 5 hours and the Yacht makes it in 2.5 hours. The steamboat is also 20 knots slower than the yacht

OpenStudy (anonymous):

From what I understand, the question wants me to write an equation to represent the problem. I have no idea what that equation would be. I dont need the answer to the equation, just the equation itself.

OpenStudy (perl):

$$\large \text{From the definition of speed, we know that}\\ \Large \color{blue}{\text { Distance = Rate * Time}}\\[0.1 in] \text {We are given:}\\ \Large d_s= r_s5\\ \Large d_y = r_y ~ 2.5\\ \Large d_s = d_y \\ \Large r_y = 20 + r_s \\ \Large \Rightarrow \\ \Large \\ \Large r_s\cdot 5 = r_y \cdot 2.5\\ \text {substitute}\\ \Large r_s\cdot 5 =(20+r_s) \cdot 2.5 $$ now you can solve this for steambot speed r_s

OpenStudy (perl):

the subscripts s , y stand for steamboat and yacht respectively